# Beam-Beam Backgrounds for the Cool Copper Collider

### v1

@@ 1 Introduction
Beam-beam backgrounds pose a fundamental constraint for proposed future $e^{+}e^{-}$ colliders. The intense electromagnetic (EM) fields of colliding bunches produce beamstrahlung photons, leading to the production of secondary particles [1, 2]. The dominant mechanisms for $e^{+}e^{-}$ colliders in the Higgs factory beam-parameter regime, ordered by magnitude of impact, are $e^{+}e^{-}$ incoherent pair creation (IPC) and photon-photon ($\gamma\gamma$) hadron photoproduction (HPP).
@@
Accurate modeling requires a multi-step simulation chain that includes beam-beam interactions, event generation, and the propagation of background particles through the detector volume, ensuring the proper estimation of the resulting occupancies and impact on detector performance. The future collider software ecosystem based on Key4hep [3] provides a modular, scalable framework for these studies, having previously been validated through end-to-end pipelines combining beam-beam generators, hadronic event generators, and Geant4 simulations for the International Linear Collider (ILC), the Compact Linear Collider (CLIC) and the Future Circular Collider (FCC) [4].
@@ 2 Cool Copper Collider and its beam parameters
C3 is a linear $e^{+}e^{-}$ collider concept that utilizes compact, high-gradient, normal-conducting accelerators with distributed coupling, operated at around 80K80\ $\mathrm${K} [5, 6]. The design of C3 follows a comprehensive optimization strategy, integrating considerations of the main linacs, collider subsystems, and beam dynamics to achieve the necessary luminosity at minimal overall cost. It is specifically tailored for a physics program at center-of-mass (CoM) energies of 250 and 550 GeV, enabling precision Higgs studies and measurements of Higgs self-coupling. The C3 beam and machine parameters are selected to match the luminosity profile of ILC [7, 8] and the proposed Linear Collider Facility (LCF) at CERN [9, 10]. The entire facility spans 8 km, sufficient to support both energy stages. The initial 250 GeV operation can be upgraded to 550 GeV by incorporating additional radio frequency (RF) power sources into the main linac. This upgrade strategy is feasible because the increased gradient and power demand are counterbalanced by modifications to the beam format, ensuring a constant beam-loading fraction and preserving RF efficiency. By refining the cavity geometry, optimizing the RF distribution, and leveraging the enhanced conductivity of copper at liquid nitrogen temperatures, the peak power demands on high-power RF sources are substantially reduced [11]. This enables a high beam-loading fraction, approaching 50% [12], resulting in a compact and efficient collider design. Through the implementation of distributed-coupling and cryogenic copper technology, peak RF power requirements are reduced by a factor of six [12, 13], comparing a cold distributed coupling structure to the effective shunt impedance of a traveling wave high-gradient linac, significantly improving overall efficiency.
@@
The C3 machine parameters have recently been optimized in [8, 14] with the purpose of delivering high luminosity at the Interaction Point (IP) without increasing site power. The current optimized parameters for C3 are listed in Table 1 and consist of a baseline scenario, which delivers the same instantaneous luminosity as ILC [7], a sustainability-update scenario that maintains luminosity while reducing site power by around $30\%$ by halving the bunch spacing and doubling the number of bunches, as well as a high-luminosity scenario.
@@
The beam parameters at the IP follow the initial Parameter Set 1 (PS1) for the baseline and sustainability-update scenarios at 250 GeV, which is summarized in Table˜2. For the other scenarios, the revised parameter set, referred to as Parameter Set 2 (PS2), is used, which achieves a $\sim 40\%$ increase in total luminosity, while ensuring that beamstrahlung-induced backgrounds remain at acceptable levels. The changes with respect to PS1 are a reduction in the vertical emittance $\epsilon_{y}^{*}$ from 20 nm to 12 nm and the introduction of a vertical waist shift ($w_{y}$) of 80 µm. Additionally, a moderate increase in the horizontal emittance ($\epsilon^{*}_{x}$) from 900 nm to 1000 nm effectively controls beam-beam interaction effects, thus maintaining a manageable beam-induced background in the detector. The feasibility of an emittance reduction has been studied through a main-linac beam dynamics analysis in [15]. More details on the optimization are given in [8].
@@
Table 1: Machine-level and beam-beam background related parameters for different C3 operating scenarios: baseline (BL), sustainability update (s.u.), and high-luminosity (high-$\mathscr{L}$) at CoM energies $\sqrt{s}=250$ and 550 GeV. The bunch charge is 1 nC in all cases, and the crossing angle is 14 mrad, compensated by crab crossing [16]. For the beam-beam related quantities, the photon-photon luminosity $\mathscr{L}_{\gamma\gamma}$ per bunch crossing (BX) is given, as well as the average number $N_{\mathrm{IPC}}$ of IPC particles produced and the average number $N_{\mathrm{HPP}}$ of HPP events per BX and for an entire bunch train. Each HPP event includes, on average, 7 particles at 250 GeV, rising to 16 particles at 550 GeV.
@@
Table 2: Main parameters at the IP for the two C3 beam-dynamics working points. Adapted from [8].
@@ 3 Simulation framework
Our simulation pipeline integrates established tools into a reproducible end-to-end chain. Beam-beam interactions are simulated using Guinea-Pig [1, 17] or its modern C++ variant Guinea-Pig++ [18], capturing non-linear pinch effects, beamstrahlung, and coherent/incoherent processes. Photon-induced hadronic backgrounds are generated by interfacing with event generators such as Whizard [19] and Pythia [20], using photon spectra obtained from Circe [21]. The detector description is performed with DD4hep [22], which is interfaced with Geant4 [23] for the full propagation of particles through the detector volume. Using the ddsim [24] toolkit, we obtain the simulated hits (SimHits) that follow the EDM4hep  [25] Event Data Model. Figure 1 summarizes the simulation workflow and the diagnostics we retain at each stage.
@@
This workflow is parameter–driven and modular. Beam parameters, such as energy, bunch size, and charge, enter through a single configuration, and the beam–beam stage may accept multiple inputs: while we use Guinea-Pig as default, alternative codes such as CAIN [26] and WARPX [27] can be accommodated with minimal adjustments. On the detector side, the pipeline works with any DD4hep geometry – k4geo models for SiD, ILD, CLD, IDEA, ALLEGRO, and other detector concepts are available [28]; hits are persisted in EDM4hep so downstream timing and occupancy studies remain identical across detector choices and beam parameter scans.
@@ 3.1 Detector simulation
For the studies presented in this work, the SiD detector concept [29], originally developed for the ILC, is used. SiD features all-silicon vertex and tracker systems, a silicon-tungsten electromagnetic calorimeter (ECAL), and a scintillator-steel hadronic calorimeter (HCAL), all enclosed within a 55\,T solenoid and muon spectrometers interleaved with the flux return steel. It also includes two forward calorimetry systems, the luminosity calorimeter (LumiCal) and the beam calorimeter (BeamCal), intended for luminosity and beam condition monitoring. The main SiD parameters are given in Appendix˜A.
For the purposes of this study, we have assumed the SiD_o2_v04 geometry description, as implemented in k4geo [30]. This description keeps geometry, materials, and sensitive surfaces in a single hierarchy and guarantees bit‑identical simulation between fast prototyping studies and the final digitisation/reconstruction chain.
@@
For the cell sizes, notably, silicon pixels with 10 µm pitch based on the Monolithic Active Pixel Sensor (MAPS) technology are assumed for the vertex, whereas for the tracker and ECAL, large-area MAPS are also considered. Simulation studies for the SiD ECAL [31] demonstrate the superior EM energy resolution and shower separation of MAPS compared to the target in the ILC Technical Design Report [32]. Significant progress has been made in the development of MAPS sensors for applications in vertex and tracking detectors, as exemplified by the ALICE detector upgrades [33, 34].
@@
Figure 2: View of the SiD detector in the rzr-z plane showing the expected number of hits from the IPC and HPP backgrounds in each subdetector system per bunch crossing, assuming the C3-250 PS1 parameter scenario.
@@ 4.1 Beamstrahlung
where $N_{e}$ is the bunch population, $\gamma$ is the Lorentz factor of the beam particles, and $\sigma_{x,y}^{*}$ and $\sigma_{z}^{*}$ are the transverse RMS bunch sizes and the bunch length, respectively. It expresses the average field strength experienced by beam particles in units of the Schwinger critical field and is a crucial parameter that must be kept at as low levels as possible in order to suppress photon production at the IP. For C3, $\langle\Upsilon\rangle$ takes values of about 0.06 at 250 GeV and 0.20 at 550 GeV, indicating that beamstrahlung is significant but still below the regime where quantum corrections dominate. The average fractional energy loss is roughly 3% at 250 GeV and increases to 9% at 550 GeV [8], while the luminosity spectrum remains sufficiently narrow as required for precision physics measurements.
@@ 4.2 Incoherent pair production
IPC is the primary beam–beam background at C3, arising from quantum electrodynamic (QED) scatterings between beamstrahlung photons and virtual photons from the opposing bunch, rather than from the collective bunch field. Note that, additional beamstrahlung-induced processes, such as trident cascades and coherent pair production, have a negligible cross-section for C3 parameters and are not further analyzed in this study [8, 35]. IPC consists of three main subprocesses, ordered by relative contribution [36, 35]: (i) Bethe–Heitler (BH), $\gamma\gamma^{\ast}\to e^{+}e^{-}$; (ii) linear Breit–Wheeler (BW), $\gamma\gamma\to e^{+}e^{-}$; and (iii) Landau–Lifshitz (LL), $\gamma^{\ast}\gamma^{\ast}\to e^{+}e^{-}$.
The IPC background is simulated using the Particle-In-Cell code Guinea-Pig++. Beams are represented by macro-particles that are longitudinally sliced and distributed onto a 2D mesh of cells; the code advances the bunches in time-steps, deposits the macro-particle charges onto the mesh, solves the fields on the cell nodes, and pushes the particles forward accordingly. During the collision, Guinea-Pig++: (i) emits beamstrahlung photons, (ii) constructs equivalent-photon fluxes, and (iii) generates $e^{+}e^{-}$ pairs from each of the three incoherent subprocesses [36, 35].
The effective cross section for the sum of the three processes, taking into account virtuality and beam size effects, is $\mathscr{O}(10-100)\ \mathrm{mb}$ [36], leading to $\mathscr{O}(10^{4}-10^{5})$ $e^{\pm}$ per bunch crossing. The exact values for each C3 scenario are given in Table˜1.
@@ 4.3 Hadron photoproduction
Beyond the dominant IPC background, photon-photon interactions at C3 generate hadronic final states through quark pair production and subsequent hadronization. Although subdominant in cross section — about 0.3 µb — HPP produces a more centrally distributed hadronic activity that can significantly impact calorimeter performance. Hadron photoproduction events are generated by using the parameterized Circe 2 photon spectra to provide initial momenta to a dedicated $\gamma\gamma\!\to$ hadrons model in Whizard 3.1.5 [37]. The implementation follows the Barklow–Chen–Peskin prescription of separating soft and jet-like components at low hadronic invariant mass, tuned with PETRA and LEP data, and then using perturbative quantum chromodynamics and a hadronization model at higher masses [38]. The diphoton CoM energy at which this transition occurs is configurable and set to $\sqrt{s_{\gamma\gamma}}=3\ \mathrm{GeV}$. Concretely, for $\sqrt{s_{\gamma\gamma}}$ values below this threshold, we use a model constrained to historical $e^{+}e^{-}$/two-photon measurements, while for larger values of $\sqrt{s_{\gamma\gamma}}$ events are showered and hadronized with Pythia 6.4 [20], preserving energy-momentum and flavor correlations across the transition [39].
The $\gamma\gamma$ luminosity spectrum at C3, as extracted from Guinea-Pig++, is parameterized in Circe 2  [21] as a function of the energy fractions $x,y$ of the colliding photons originating from the electron and positron beams, respectively. This approach captures the full spectrum of virtual and real photon interactions, including beamstrahlung-enhanced contributions at low photon energies. The differential photon luminosity as a function of the photon-photon CoM energy, normalized to the nominal beam center-of-mass energy $\sqrt{s_{\gamma\gamma}}/\sqrt{s}$, is shown in Fig.˜3 for different C3 beam parameter sets at both $\sqrt{s}=$ 250 and 550 GeV.

### v2

@@ 1 Introduction
Beam-beam backgrounds pose a fundamental constraint for proposed future $e^{+}e^{-}$ colliders. The intense electromagnetic (EM) fields of colliding bunches produce beamstrahlung photons, leading to the production of secondary particles [Schulte:1996, Yokoya:1991qz]. The dominant mechanisms for $e^{+}e^{-}$ colliders in the Higgs factory beam-parameter regime, ordered by magnitude of impact, are $e^{+}e^{-}$ incoherent pair creation (IPC) and photon-photon ($\gamma\gamma$) hadron photoproduction (HPP).
@@
Accurate modeling requires a multi-step simulation chain that includes beam-beam interactions, event generation, and the propagation of background particles through the detector volume, ensuring the proper estimation of the resulting occupancies and impact on detector performance. The future collider software ecosystem based on Key4hep [Sailer:2020fah] provides a modular, scalable framework for these studies, having previously been validated through end-to-end pipelines combining beam-beam generators, hadronic event generators, and Geant4 simulations for the International Linear Collider (ILC), the Compact Linear Collider (CLIC) and the Future Circular Collider (FCC) [Ganis:2021vgv].
@@ 2 Cool Copper Collider and its beam parameters
C3 is a linear $e^{+}e^{-}$ collider concept that utilizes compact, high-gradient, normal-conducting accelerators with distributed coupling, operated at around 80K80\ $\mathrm${K} [vernieri2023cool, dasu2022strategy]. The design of C3 follows a comprehensive optimization strategy, integrating considerations of the main linacs, collider subsystems, and beam dynamics to achieve the necessary luminosity at minimal overall cost. It is specifically tailored for a physics program at center-of-mass (CoM) energies of 250 and 550 GeV, enabling precision Higgs studies and measurements of Higgs self-coupling. The C3 beam and machine parameters are selected to match the luminosity profile of ILC [ILC_Snowmass, c3_lumi] and the proposed Linear Collider Facility (LCF) at CERN [LinearColliderVision:2025hlt, LinearCollider:2025lya]. The entire facility spans 8 km, sufficient to support both energy stages. The initial 250 GeV operation can be upgraded to 550 GeV by incorporating additional radio frequency (RF) power sources into the main linac. This upgrade strategy is feasible because the increased gradient and power demand are counterbalanced by modifications to the beam format, ensuring a constant beam-loading fraction and preserving RF efficiency. By refining the cavity geometry, optimizing the RF distribution, and leveraging the enhanced conductivity of copper at liquid nitrogen temperatures, the peak power demands on high-power RF sources are substantially reduced [nanni2023status]. This enables a high beam-loading fraction, approaching 50% [Bane:2018fzj], resulting in a compact and efficient collider design. Through the implementation of distributed-coupling and cryogenic copper technology, peak RF power requirements are reduced by a factor of six [Bane:2018fzj, Grudiev2010DesignOT], comparing a cold distributed coupling structure to the effective shunt impedance of a traveling wave high-gradient linac, significantly improving overall efficiency.
@@
The C3 machine parameters have recently been optimized in [c3_lumi, C3_ESPPU] with the purpose of delivering high luminosity at the Interaction Point (IP) without increasing site power. The current optimized parameters for C3 are listed in Table 1 and consist of a baseline scenario, which delivers the same instantaneous luminosity as ILC [ILC_Snowmass], a sustainability-update scenario that maintains luminosity while reducing site power by around $30\%$ by halving the bunch spacing and doubling the number of bunches, as well as a high-luminosity scenario.
@@
The beam parameters at the IP follow the initial Parameter Set 1 (PS1) for the baseline and sustainability-update scenarios at 250 GeV, which is summarized in Table˜2. For the other scenarios, the revised parameter set, referred to as Parameter Set 2 (PS2), is used, which achieves a $\sim 40\%$ increase in total luminosity, while ensuring that beamstrahlung-induced backgrounds remain at acceptable levels. The changes with respect to PS1 are a reduction in the vertical emittance $\epsilon_{y}^{*}$ from 20 nm to 12 nm and the introduction of a vertical waist shift ($w_{y}$) of 80 µm. Additionally, a moderate increase in the horizontal emittance ($\epsilon^{*}_{x}$) from 900 nm to 1000 nm effectively controls beam-beam interaction effects, thus maintaining a manageable beam-induced background in the detector. The feasibility of an emittance reduction has been studied through a main-linac beam dynamics analysis in [Tan:2025rss]. More details on the optimization are given in [c3_lumi].
@@
Table 1: Machine-level and beam-beam background related parameters for different C3 operating scenarios: baseline (BL), sustainability update (s.u.), and high-luminosity (high-$\mathscr{L}$) at CoM energies $\sqrt{s}=250$ and 550 GeV. The bunch charge is 1 nC in all cases, and the crossing angle is 14 mrad, compensated by crab crossing [adolphsen2013internationallinearcollidertechnical]. For the beam-beam related quantities, the photon-photon luminosity $\mathscr{L}_{\gamma\gamma}$ per bunch crossing (BX) is given, as well as the average number $N_{\mathrm{IPC}}$ of IPC particles produced and the average number $N_{\mathrm{HPP}}$ of HPP events per BX and for an entire bunch train. Each HPP event includes, on average, 7 particles at 250 GeV, rising to 16 particles at 550 GeV.
@@
Table 2: Main parameters at the IP for the two C3 beam-dynamics working points. Adapted from [c3_lumi].
@@ 3 Simulation framework
Our simulation pipeline integrates established tools into a reproducible end-to-end chain. Beam-beam interactions are simulated using Guinea-Pig [Schulte:1996, Schulte:1999] or its modern C++ variant Guinea-Pig++ [Rimbault:2007wfy], capturing non-linear pinch effects, beamstrahlung, and coherent/incoherent processes. Photon-induced hadronic backgrounds are generated by interfacing with event generators such as Whizard [Kilian:2007gr] and Pythia [Sj_strand_2006], using photon spectra obtained from Circe [Ohl:1996fi]. The detector description is performed with DD4hep [Frank:2014zya], which is interfaced with Geant4 [Agostinelli:2002hh] for the full propagation of particles through the detector volume. Using the ddsim [Petric:2017psf] toolkit, we obtain the simulated hits (SimHits) that follow the EDM4hep  [Gaede:2021izq] Event Data Model. Figure 1 summarizes the simulation workflow and the diagnostics we retain at each stage.
@@
This workflow is parameter–driven and modular. Beam parameters, such as energy, bunch size, and charge, enter through a single configuration, and the beam–beam stage may accept multiple inputs: while we use Guinea-Pig as default, alternative codes such as CAIN [CAIN] and WARPX [warpx] can be accommodated with minimal adjustments. On the detector side, the pipeline works with any DD4hep geometry – k4geo models for SiD, ILD, CLD, IDEA, ALLEGRO, and other detector concepts are available [k4geo]; hits are persisted in EDM4hep so downstream timing and occupancy studies remain identical across detector choices and beam parameter scans.
@@ 3.1 Detector simulation
For the studies presented in this work, the SiD detector concept [Aihara:2009ad], originally developed for the ILC, is used. SiD features all-silicon vertex and tracker systems, a silicon-tungsten electromagnetic calorimeter (ECAL), and a scintillator-steel hadronic calorimeter (HCAL), all enclosed within a 55\,T solenoid and muon spectrometers interleaved with the flux return steel. It also includes two forward calorimetry systems, the luminosity calorimeter (LumiCal) and the beam calorimeter (BeamCal), intended for luminosity and beam condition monitoring. The main SiD parameters are given in Appendix˜A.
For the purposes of this study, we have assumed the SiD_o2_v04 geometry description, as implemented in k4geo [k4geo_sid_o2_v04]. This description keeps geometry, materials, and sensitive surfaces in a single hierarchy and guarantees bit‑identical simulation between fast prototyping studies and the final digitization/reconstruction chain. The machine-detector interface (MDI) region, including the beampipe and forward shielding, is an integral part of the k4geo geometry description and is simulated together with all other detector components. The beampipe geometry follows the SiD baseline design [behnke2013internationallinearcollidertechnical] and consists of a thin-walled central beryllium cylinder that transitions to conical sections extending toward the forward calorimetry. The central section, with an inner radius of 12 mm, is positioned to clear the IPC background envelopes shown in Fig.˜5.
@@
For the cell sizes, notably, silicon pixels with 10 µm pitch based on the Monolithic Active Pixel Sensor (MAPS) technology are assumed for the vertex, whereas for the tracker and ECAL, large-area MAPS are also considered. Simulation studies for the SiD ECAL [Brau:2022sxr] demonstrate the superior EM energy resolution and shower separation of MAPS compared to the target in the ILC Technical Design Report [behnke2013internationallinearcollidertechnical]. Significant progress has been made in the development of MAPS sensors for applications in vertex and tracking detectors, as exemplified by the ALICE detector upgrades [ALICE:2023udb, Groettvik:2024onw].
@@
Figure 2: Quarter view of the SiD detector in the rzr-z plane showing the expected number of hits from the IPC and HPP backgrounds in each subdetector system per bunch crossing, assuming the C3-250 PS1 parameter scenario. The inset provides an enlarged view of the vertex region.
@@ 4.1 Beamstrahlung
where $N_{e}$ is the bunch population, $\gamma$ is the Lorentz factor of the beam particles, and $\sigma_{x,y}^{*}$ and $\sigma_{z}^{*}$ are the transverse RMS bunch sizes and the bunch length, respectively. It expresses the average field strength experienced by beam particles in units of the Schwinger critical field and is a crucial parameter that must be kept at as low levels as possible in order to suppress photon production at the IP. For C3, $\langle\Upsilon\rangle$ takes values of about 0.06 at 250 GeV and 0.20 at 550 GeV, indicating that beamstrahlung is significant but still below the regime where quantum corrections dominate. The average fractional energy loss is roughly 3% at 250 GeV and increases to 9% at 550 GeV [c3_lumi], while the luminosity spectrum remains sufficiently narrow as required for precision physics measurements.
@@ 4.2 Incoherent pair production
IPC is the primary beam–beam background at C3, arising from quantum electrodynamic (QED) scatterings between beamstrahlung photons and virtual photons from the opposing bunch, rather than from the collective bunch field. Note that, additional beamstrahlung-induced processes, such as trident cascades and coherent pair production, have a negligible cross-section for C3 parameters and are not further analyzed in this study [c3_lumi, Esberg:2014]. IPC consists of three main subprocesses, ordered by relative contribution [Rimbault:2006, Esberg:2014]: (i) Bethe–Heitler (BH), $\gamma\gamma^{\ast}\to e^{+}e^{-}$; (ii) linear Breit–Wheeler (BW), $\gamma\gamma\to e^{+}e^{-}$; and (iii) Landau–Lifshitz (LL), $\gamma^{\ast}\gamma^{\ast}\to e^{+}e^{-}$.
The IPC background is simulated using the Particle-In-Cell code Guinea-Pig++. Beams are represented by macro-particles that are longitudinally sliced and distributed onto a 2D mesh of cells; the code advances the bunches in time-steps, deposits the macro-particle charges onto the mesh, solves the fields on the cell nodes, and pushes the particles forward accordingly. During the collision, Guinea-Pig++: (i) emits beamstrahlung photons, (ii) constructs equivalent-photon fluxes, and (iii) generates $e^{+}e^{-}$ pairs from each of the three incoherent subprocesses [Rimbault:2006, Esberg:2014].
The effective cross section for the sum of the three processes, taking into account virtuality and beam size effects, is $\mathscr{O}(10-100)\ \mathrm{mb}$ [Rimbault:2006], leading to $\mathscr{O}(10^{4}-10^{5})$ $e^{\pm}$ per bunch crossing. The exact values for each C3 scenario are given in Table˜1.
@@ 4.3 Hadron photoproduction
Beyond the dominant IPC background, photon-photon interactions at C3 generate hadronic final states through quark pair production and subsequent hadronization. Although subdominant in cross section — about 0.3 µb — HPP produces a more centrally distributed hadronic activity that can significantly impact calorimeter performance. Hadron photoproduction events are generated by using the parameterized Circe 2 photon spectra to provide initial momenta to a dedicated $\gamma\gamma\!\to$ hadrons model in Whizard 3.1.5 [Stienemeier:2021]. The implementation follows the Barklow–Chen–Peskin prescription of separating soft and jet-like components at low hadronic invariant mass, tuned with PETRA and LEP data, and then using perturbative quantum chromodynamics and a hadronization model at higher masses [Chen:1993]. The diphoton CoM energy at which this transition occurs is configurable and set to $\sqrt{s_{\gamma\gamma}}=3\ \mathrm{GeV}$. Concretely, for $\sqrt{s_{\gamma\gamma}}$ values below this threshold, we use a model constrained to historical $e^{+}e^{-}$/two-photon measurements, while for larger values of $\sqrt{s_{\gamma\gamma}}$ events are showered and hadronized with Pythia 6.4 [Sj_strand_2006], preserving energy-momentum and flavor correlations across the transition [Bierlich:2022].
The $\gamma\gamma$ luminosity spectrum at C3, as extracted from Guinea-Pig++, is parameterized in Circe 2  [Ohl:1996fi] as a function of the energy fractions $x,y$ of the colliding photons originating from the electron and positron beams, respectively. This approach captures the full spectrum of virtual and real photon interactions, including beamstrahlung-enhanced contributions at low photon energies. The differential photon luminosity as a function of the photon-photon CoM energy, normalized to the nominal beam center-of-mass energy $\sqrt{s_{\gamma\gamma}}/\sqrt{s}$, is shown in Fig.˜3 for different C3 beam parameter sets at both $\sqrt{s}=$ 250 and 550 GeV.